OpenStudy (anonymous):

The ordered pairs (1, 1), (2, 4), (3, 9), (4, 16), and (5, 25) represent a function. What is a rule that represents this function? A. y = x2 B. y = 2x C. y = x + 2 D. y = 2x

OpenStudy (whpalmer4):

Try the various functions out and see if they produce the data. You'll need to try several points, as some of the functions will work for some, but not all of the data points. Remember, ordered pairs are interpreted as \((x,y)\) so \((2,4)\) means \(x = 2, y = 4\) As an example: let's try answer choice C: \(y = x+2\) First data point I'll try is \((2,4)\): \[y = x+2\]\[4 = 2+2\]\[4=4\checkmark\]that data point works, so answer C might be correct. Let's try another data point, \((1,1)\) \[y=x+2\]\[1 = 1+2\]\[1=3\]No, that is not true, so the function \(y = x+2\) does not correctly produce all of the data. Answer choice C is not correct.

OpenStudy (anonymous):

thanks!!